OpenStudy (dtan5457):

If (x+iy)^2=15+8i, and x and y are real #'s. and x>0, what is x-y?

zepdrix (zepdrix):

Expanding out the left side gives us:\[\large\rm x^2+2xyi-y^2\quad=\quad 15+8i\]

zepdrix (zepdrix):

So we have:\[\large\rm \color{orangered}{x^2-y^2}+\color{royalblue}{2xyi}\quad=\quad \color{orangered}{15}+\color{royalblue}{8i}\]The real parts have to match up while the imaginary parts match up as well, ya?

zepdrix (zepdrix):

\[\large\rm x^2-y^2=15\]Applying difference of squares,\[\large\rm (x-y)(x+y)=15\]And we have our other equation as well,\[\large\rm 2xy=8\]

zepdrix (zepdrix):

So what can we do with that information... hmmm

OpenStudy (dtan5457):

what happened to the i again?

zepdrix (zepdrix):

Matching up the imaginary parts gives us this,\[\large\rm 2xyi=8i\qquad\implies\qquad 2xy=8\]

OpenStudy (dtan5457):

u can just equal them like that?

zepdrix (zepdrix):

Yes :) If you have an equivalence like this:\[\large\rm \color{orangered}{x^2-y^2}+\color{royalblue}{2xyi}\quad=\quad \color{orangered}{15}+\color{royalblue}{8i}\]The real on the left must be equal to the real on the right.\[\large\rm \color{orangered}{x^2-y^2=15}\]And the imaginary on the left must be equal to the imaginary on the right.\[\large\rm \color{royalblue}{2xyi=8i\qquad\to\qquad 2xy=8}\]

zepdrix (zepdrix):

But I'm getting stuck after that >.< Hmm thinking...

OpenStudy (dtan5457):

what if we don't remove any i's, and just do \[(x-y)=\frac{ 15+8i }{ x+y }-2xyi\]

zepdrix (zepdrix):

@dan815 @Kainui @Michele_Laino @phi Hmm :d

OpenStudy (dtan5457):

@amistre64

OpenStudy (amistre64):

zeps idea is more sound to me

OpenStudy (amistre64):

x-y = 15/(x+y) and y=4/x

OpenStudy (amistre64):

does x-y have a single solution?

OpenStudy (dtan5457):

it does not say

OpenStudy (amistre64):

the hint that x > 0 seems to imply (x-y) = 15/(x+4/x)

OpenStudy (amistre64):

well, x=4 and y=1 seems to suffice for the constraints, giving x-y = 3

OpenStudy (dtan5457):

oh yeah that does work. still a bit confused on the work though

OpenStudy (amistre64):

zep did most of the process ... if 2 things are equal then their parts must be equal

OpenStudy (phi):

we could solve the quartic, but I'm guessing there is a tricky way to do this quickly

OpenStudy (amistre64):

(x^2-y^2) + 2xy i =========== 15 + 8i

OpenStudy (phi):

\[ x^2 -y^2 = 15 \\ y= \frac{4}{x} \\ x^2 - \frac{16}{x^2} -15 =0 \] \[ x^4 -15 x^2 -16= 0 \\ u^2 -15u -16=0 \] (u-16)(u+1)= 0 u=16 x^2 =16 x=4 y=1 x-y=3

OpenStudy (dtan5457):

is it possible to solve for the term (x+y) in this problem?

OpenStudy (phi):

5 would be one solution

OpenStudy (dtan5457):

but say I had an approach of (x+y)(x-y)=15+8i-2xyi xy=4 (x+y)(x-y)=15 could i get x+y just knowing what i have so far?

OpenStudy (phi):

I only know the way I did it. If we know x and y are real i.e. not imaginary or complex, and x>0, then we would find x=4 , y=1 and x+y= 5 if there is a nicer way to do it, I would like to see it

OpenStudy (dtan5457):

alright, my teacher will go over this tomorrow so i'll see which method he decides to use. thank you both

OpenStudy (phi):

if you get any insight, please post it here!